Mole fraction of solute in aqueous solution of 30% NaOH
 

Grade:12th pass

1 Answers

aparazita sharma
5 years ago
let mass of solution be 100 g then mass of naoh is 30 g and mass of water is 100-30= 70 g
now moles of naoh = mass of naoh / molar mass of naoh 
 =  30 / 40 = 0.75
similarly moles of water = 70 /18 =3.88
so total moles=0.75+3.88=4.63
mole fraction of solute = moles of naoh / total moles
                                        0.75/ 4.63=0.16 or 27/ 167